one number is 7 less than another number. if 5 times the smaller number is subtracted from twicw the large number, the result is -10. find the numbers.
Answer by Paul(988)

x-7=y
y=-7+x
5x-2y=-10
5x-2(7+x)=-10
5x+14-x=-10
4x=24
x=-6